OpenStudy (anonymous):

What is the volume of a cone (in cubic inches) with a radius of 2 inches and a height of 3 inches? 12.56 cubic inches 25.12 cubic inches 37.68 cubic inches 56.52 cubic inches

OpenStudy (anonymous):

The formula is \(\dfrac{1}{3}\) x r x h, where r is the radius, and h is the height.

OpenStudy (anonymous):

how do you do it though

OpenStudy (anonymous):

So can you figure out \(\dfrac{1}{3}\) x 2 x 3?

OpenStudy (campbell_st):

Actually the volume formula is \[V = \frac{1}{3} \pi r^2 h\] so substitute the values r = 2 and h = 3 the calculate the answer

OpenStudy (anonymous):

so is it 18?

OpenStudy (anonymous):

Thanks. @campbell_st

OpenStudy (anonymous):

is it 18?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

First, what is 2^2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4 x 3

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Then what is \(\dfrac{1}{3}\) x 12?

OpenStudy (anonymous):

144

OpenStudy (anonymous):

?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

36

OpenStudy (anonymous):

1/3 is a fraction..

OpenStudy (anonymous):

i know but if you divide 1/3 it is 3

OpenStudy (anonymous):

1/3 is about 0.333

OpenStudy (anonymous):

No, it's 0.333

OpenStudy (anonymous):

Approximately

OpenStudy (anonymous):

ok but i think that is wrong

OpenStudy (anonymous):

im not going to medal you until you tell me the right way!

OpenStudy (campbell_st):

ok.... here is what you need to know